# Field-induced states and thermodynamics of the frustrated Heisenberg antiferromagnet on a square lattice

### v1

@@ I Introduction
The $J_{1}$$J_{2}$ square-lattice model was introduced in the late 1980s with the aim to describe the breakdown upon doping of Néel antiferromagnetic (NAF) state in the high-temperature cuprate superconductors [3, 4, 5]. Over the last few decades the zero-field case, $h=0$, has attracted much attention, since it is a canonical model to study frustration-induced quantum phase transitions between semiclassical ground-state phases with a magnetic long-range order (LRO) and disordered quantum phases, see, e.g., Refs. [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59].
Based on numerous studies using a plethora of quantum many-body methods, consensus has been achieved that the $J_{1}$-$J_{2}$ model in the extreme quantum limit $s=1/2$ exhibits a magnetically disordered spin-rotation-invariant quantum phase in a small region $J_{2}^{c1}<J_{2}<J_{2}^{c2}$ around $J_{2}=0.5J_{1}$. The semi-classical ground-state phases of the model, namely the Néel antiferomagnetic LRO at small $J_{2}/J_{1}<J_{2}^{c1}$ and the collinear antiferromagnetic LRO at large $J_{2}/J_{1}>J_{2}^{c2}$, are well understood. Concerning the critical values $J_{2}^{c1}$ and $J_{2}^{c2}$, some estimates are $J_{2}^{c1}=0.44-0.46$ and $J_{2}^{c2}=0.59-0.5$ [20, 32, 36, 37, 40, 44, 47, 59]. However, some recent PEPS approaches [38, 42, 41, 52] lead to a transition point $J_{2}^{c1}$ close to $0.5$ or even above. Concerning the nature of the intermediate magnetically disordered quantum phase, there is an ongoing active controversial debate [17, 20, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 44, 45, 46, 48, 50, 52, 58]. A particular controversy concerns the very existence of an excitation gap in the intermediate quantum phase. A finite excitation gap was reported in Refs. [23, 27, 28, 33], whereas indications of a gapless spin liquid state were found in Refs. [29, 30, 36, 50]. Investigations based on density matrix renormalization group with explicit implementation of $SU(2)$ spin rotation symmetry in Ref. [32] report on a gapless spin liquid for $0.44<J_{2}/J_{1}<0.5$ and a gapped plaquette valence bond phase for $0.5<J_{2}/J_{1}<0.61$. This structure of the phase diagram is reaffirmed in Ref. [50], but the point separating the two intermediate phases is shifted to $J_{2}/J_{1}\approx 0.56$. On the other hand, a recent matrix-product state method claims that there is no spin-liquid phase in this model [52].
@@
The unconventional quantum nature of the zero-field ground state in the strong-frustration regime has also stimulated the search for unconventional features in the magnetization curve [71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82] as well as finite-temperature properties [83, 84, 85, 86, 87, 88, 89, 90, 91] of the spin-1/2 FSAFM. In particular, in the regime of strong frustration $J_{2}/J_{1}\approx 0.5$, where the zero-field quantum ground state does not exhibit magnetic LRO, evidence for a well-pronounced plateau at $1/2$ of saturation was found [72].
@@ II.1 Ground states
We begin with the classical $J_{1}$-$J_{2}$ model (1) treating spins $\bm{s}_{i}$ as three-component unit vectors. For a special ratio of the two exchange constants $J_{2}/J_{1}=1/2$, magnetic frustration is enhanced featuring an infinite classical degeneracy for all magnetic fields below the saturation [72]. In order to see this we define a total spin of a square plaquette $\alpha$:
@@
The constraint (4) imposes three conditions on the eight angle variables for the four spins of each block. One more angle variable is eliminated by the broken $O(2)$ rotation symmetry about the field direction. Overall, this counting yields four continuous variables to describe various classical ground states for a single block. Analysis of the effect of fluctuations, thermal or quantum, on a state selection can be performed using the concept of an effective biquadratic exchange, see, for example, Ref. [92]. Figure 1 sketches the relevant spin configurations that appear for four site blocks in various magnetic fields [72, 93, 94]. The “up-up-up-down” (uuuduuud) configuration has the lowest energy for magnetic fields around $h_{\rm sat}/2$ and produces the magnetization plateau with $m=1/2$. At higher fields away from the 1/2 plateau, spins tilt retaining the 3:1 splitting. For fields below $m<1/2$, the noncollinear spin configuration has a more complex 2:1:1 structure. The 2:2 state, which is one of the classical ground states in the entire range $0<h<h_{\rm sat}$, is stable, in the presence of a biquadratic exchange, in the low-field region only [94].
Since the plaquettes $\alpha$ share edges, one spin participates in several plaquettes. As a result, one faces the problem of extending a single block solution to the entire lattice. This leads to additional degeneracy for the lattice spin model for each of the single block solutions in Fig. 1. The corresponding degeneracy is most easily visualized for the uuuduuud state [72]. Specifically, consider the most symmetric uuuduuud configuration. For that we choose the down spin, for example, in the lowest left corner of the first square plaquette and extend this configuration in a translationally invariant way over the whole lattice. This state consists of parallel, vertical and horizontal, lines of up-down spins separated by lines of up spins. The alternating up-down spins along any of these lines can be freely rotated around an arbitrary angle without changing the total classical energy [72]. Since the uuuduuud state permits a maximum of such deformations, one may expect it to be selected by thermal fluctuations via the “order-by-disorder” mechanism [95, 96, 97, 98, 99, 100]. A subtle property of the order-by-disorder mechanism is a competition between a local fluctuation contribution manifested by the effective biquadratic coupling and the extended zero-energy modes related to spin distribution over the entire lattice. Furthermore, in the latter case there is the possibility of distinct effects from the thermal fluctuations, which rely on the entropic contribution of the low-energy (zero) modes, and from the quantum fluctuations, with magnons at different momenta contributing equally. As we shall see below, all these exotic features of the order-by-disorder effect are relevant for the FSAFM.
@@ II.2 Field-temperature phase diagram
Figure 2 presents the constructed $h$-$T$ phase diagram for $J_{2}/J_{1}=1/2$. First, we observe the expected stabilization of the uuuduuud phase by thermal fluctuations in the regime $T\lesssim J_{1}/10$. The stability region is asymmetric around $h_{c}=0.5h_{\rm sat}$ having a significant extent into the low-field part of the diagram. Second, there are sizable regions with the 2:1:1 and 3:1 states just below or above the uuuduuud phase, respectively. The two noncollinear states possess the longitudinal LRO in the $s^{z}$ components accompanied by the quasi-LRO in transverse spin components. Mapping this from magnetic to bosonic language [104], these two phases can be identified as supersolids, whereas the parent uuuduuud phase is an insulating state with the density-wave crystal.
Figure 3: The field derivative of the magnetization for the classical FSAFM with $J_{2}=0.5$, $J_{1}=1$ at three different temperatures obtained on lattices with $L=32$.
@@
Figures 3 and 4 illustrate some of the data used in the construction of the phase diagram Fig. 2. First, let us consider the field derivative of the magnetization ${\rm d}m/{\rm d}h$. According to the analysis at the beginning of this section, at zero temperature $m(h)$ has a constant slope and ${\rm d}m/{\rm d}h$ remains flat for all fields below the saturation $h_{\rm sat}$, i.e., ${\rm d}m/{\rm d}h=J_{1}/8$ at $T=0$ and for $h<h_{\rm sat}$. At $T>0$, we observe in Fig. 3 the emergence of small but sharp and distinct maxima, which may be taken as first indications of field-induced phase transitions. For $h\lesssim 4\,J_{1}=h_{\rm sat}/2$, we observe a drop in ${\rm d}m/{\rm d}h$. The slope of the magnetization curve $m(h)$ remains finite and thus one does not find a strict plateau in the classical magnetization curve. Nevertheless, the reduced slope of the magnetization curve can be considered as a distinctive feature of the uuuduuud phase.
@@ III Quantum model
$\left.\omega(k_{x},k_{y})\right|_{J_{1}=J_{2}/2}$ $=$ $2sJ_{1}\left(\left(1+\cos k_{x}\right)\left(1+\cos k_{y}\right)-4\right)$ (8)
@@
$\left.\omega(k_{x},\pi)\right|_{J_{1}=J_{2}/2}=\left.\omega(\pi,k_{y})\right|_{J_{1}=J_{2}/2}=-8sJ_{1}+h\,$, (9)
@@ III.1.1 Magnetization curves
Figure 8 presents zero-temperature magnetization curves, where we normalizated the saturation magnetization to $m=1$. For reference, we include the plain square-lattice case with $J_{2}=0$ (Fig. 8(a)) that has been investigated in detail previously, see for example Ref. [111], including the $N=40$ results. Here, we find a smooth curve in the thermodynamic limit with an upwards curvature that ends in a logarithmic singularity. This behavior is characteristic of a two-dimensional quantum spin system, and has been analytically predicted for the present case by second-order spin-wave theory [112].
@@
We note that a Chern-Simons theory predicted a distinct $m=1/3$ plateau in addition to the $m=1/2$ one [76]. The $N=36$ and $48$ systems indeed exhibit a broader step for $J_{2}/J_{1}=0.55$ (Fig. 8(c)) and $0.6$ (Fig. 8(d)). However, the magnetization curves for the other system sizes appear to exhibit a smooth behavior in this region, the mean-field theory underlying the prediction of Ref. [76] is not well controlled, and the DMRG investigation Ref. [81] did not observe an $m=1/3$ plateau either. We have therefore not drawn an $m=1/3$ plateau in our extrapolated curves.
Lastly, we note that at $J_{2}/J_{1}=0.5$ and $J_{2}/J_{1}=0.55$ we have drawn an $m=0$ plateau in Fig. 8(b) and (c), respectively. This corresponds to a spin gap in zero field, as is to be expected for the paramagnetic phases around $J_{2}/J_{1}=1/2$ mentioned in the Introduction. The situation for $J_{2}/J_{1}=0.6$ (Fig. 8(d)) is less clear. Here, the $N=36$ system suggests the presence of a spin gap, but this is not confirmed by the $N=40$ system. We therefore have not drawn am $m=0$ plateau although we cannot exclude the existence of a small $m=0$ plateau, i.e., spin gap for $J_{2}/J_{1}=0.6$. In any case, we recall that the zero-field behavior is a challenging issue that is beyond the scope of the present work.
@@ III.1.2 $m=1/2$ plateau
Now we focus on a more detailed characterization of the $m=1/2$ plateau. First, we look at the width of the $m=1/2$ step, calculated as $\Delta h=E(N/2+1)-2E(N/2)+E(N/2-1)$, where $E(S^{z})$ is the energy of the ground state for the given value of $S^{z}$ at $h=0$. Figure 9 presents the behavior of $\Delta h$ as a function of $J_{2}/J_{1}$ for systems with $32\leq N\leq 52$ spins 1/2. We have chosen lattices with an aspect ratios as close as possible to unity for a given size $N$, but nevertheless the results fall into two groups, namely the square respectively rectangular geometries for $N=36$ and $48$ on the one side, and the tilted cases $N=32$, $40$, and $52$ on the other side. This non-monotonic finite-size behavior renders an extrapolation to $N\to\infty$ difficult. Nevertheless, we observe a clear and consistent enhancement of the $m=1/2$ step size in the region $0.52\lesssim J_{2}/J_{1}\lesssim 0.7$, which we thus take as a first estimate for the stability region of the $m=1/2$ plateau.
To gain further insight into the state of the $m=1/2$ plateau, Fig. 2 shows real-space spin-spin correlations at distance $R$ for the $N=52$ lattice of Fig. 7, i.e., at $S^{z}=13$. Note that $\langle s^{z}_{R}\rangle=m/2=1/4$. We therefore subtract a background of $\langle s^{z}_{0}\rangle\langle s^{z}_{R}\rangle=1/16$ from the longitudinal correlation function $\langle s^{z}_{0}s^{z}_{R}\rangle$ to obtain the connected one, that is shown in Fig. 2(a). The transverse one in Fig. 2(b) is obtained as $\langle s^{\perp}_{0}s^{\perp}_{R}\rangle=\langle\bm{s}_{0}\cdot\bm{s}_{R}\rangle-\langle s^{z}_{0}s^{z}_{R}\rangle$. Overall, we observe three distinct regimes in Fig. 2: (i) For $J_{2}/J_{1}\lesssim 0.5$ the transverse correlations dominate. This is consistent with Néel order in the transverse components, as expected in the limit $J_{2}=0$. (ii) For $0.52\lesssim J_{2}/J_{1}<0.7$, the transverse correlations essentially vanish, but strong longitudinal ones develop. This is consistent with the collinear uuuduuud state that is expected for an $m=1/2$ plateau from the investigation of the classical model in Sec. II.1. (iii) For $J_{2}/J_{1}>0.7$, the longitudinal components vanish again while a new signal develops in the transverse components. Different signs reflect an order that differs from the one in regime (i). Indeed, for $J_{1}=0$, one has two interpenetrating square lattices, each of which develops independent Néel order. A weak coupling $J_{1}>0$ then leads to a “stripe” phase, consistent with the correlations observed in Fig. 2(b) for $J_{2}/J_{1}>0.7$.

### v2

@@ I Introduction
The $J_{1}$$J_{2}$ square-lattice model was introduced in the late 1980s with the aim to describe the breakdown upon doping of the Néel antiferromagnetic (NAF) state in the high-temperature cuprate superconductors [3, 4, 5]. Over the last few decades the zero-field case, $h=0$, has attracted much attention, since it is a canonical model to study frustration-induced quantum phase transitions between semiclassical ground-state phases with a magnetic long-range order (LRO) and disordered quantum phases, see, e.g., Refs. [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59].
Based on numerous studies using a plethora of quantum many-body methods, consensus has been achieved that the $J_{1}$-$J_{2}$ model in the extreme quantum limit $s=1/2$ exhibits a magnetically disordered spin-rotation-invariant quantum phase in a small region $J_{2}^{c1}<J_{2}<J_{2}^{c2}$ around $J_{2}=0.5J_{1}$. The semi-classical ground-state phases of the model, namely the Néel antiferomagnetic LRO at small $J_{2}/J_{1}<J_{2}^{c1}$ and the collinear antiferromagnetic LRO at large $J_{2}/J_{1}>J_{2}^{c2}$, are well understood. Concerning the critical values $J_{2}^{c1}$ and $J_{2}^{c2}$, some estimates are $J_{2}^{c1}=0.44-0.46$ and $J_{2}^{c2}=0.59-0.6$ [20, 32, 36, 37, 40, 44, 47, 59]. However, some recent PEPS approaches [38, 42, 41, 52] lead to a transition point $J_{2}^{c1}$ close to $0.5$ or even above. Concerning the nature of the intermediate magnetically disordered quantum phase, there is an ongoing active controversial debate [17, 20, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 44, 45, 46, 48, 50, 52, 58]. A particular controversy concerns the very existence of an excitation gap in the intermediate quantum phase. A finite excitation gap was reported in Refs. [23, 27, 28, 33], whereas indications of a gapless spin liquid state were found in Refs. [29, 30, 36, 50]. Investigations based on density matrix renormalization group with explicit implementation of $SU(2)$ spin rotation symmetry in Ref. [32] report on a gapless spin liquid for $0.44<J_{2}/J_{1}<0.5$ and a gapped plaquette valence bond phase for $0.5<J_{2}/J_{1}<0.61$. This structure of the phase diagram is reaffirmed in Ref. [50], but the point separating the two intermediate phases is shifted to $J_{2}/J_{1}\approx 0.56$. On the other hand, a recent matrix-product state method claims that there is no spin-liquid phase in this model [52].
@@
The unconventional quantum nature of the zero-field ground state in the strong-frustration regime has also stimulated the search for unconventional features in the magnetization curve [71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83] as well as finite-temperature properties [16, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93] of the spin-1/2 FSAFM. In particular, in the regime of strong frustration $J_{2}/J_{1}\approx 0.5$, where the zero-field quantum ground state does not exhibit magnetic LRO, evidence for a well-pronounced plateau at $1/2$ of saturation was found [72].
@@ II.1 Ground states
We begin with the classical $J_{1}$-$J_{2}$ model (1) treating spins $\bm{s}_{i}$ as three-component unit vectors. For the special ratio of the two exchange constants $J_{2}/J_{1}=1/2$, magnetic frustration is enhanced featuring an infinite classical degeneracy for all magnetic fields below the saturation [72]. In order to see this we define the total spin of a square plaquette $\alpha$:
@@
The constraint (4) imposes three conditions on the eight angle variables for the four spins of each block. One more angle variable is eliminated by the broken $O(2)$ rotation symmetry about the field direction. Overall, this counting yields four continuous variables to describe various classical ground states for a single block. Analysis of the effect of fluctuations, thermal or quantum, on a state selection can be performed using the concept of an effective biquadratic exchange, see, for example, Ref. [94]. Figure 1 sketches the relevant spin configurations that appear for four-site blocks in various magnetic fields [72, 95, 96]. The “up-up-up-down” (uuuduuud) configuration has the lowest energy for magnetic fields around $h_{\rm sat}/2$ and produces the magnetization plateau with $m=1/2$. At higher fields away from the 1/2 plateau, spins tilt retaining the 3:1 splitting. For fields below $m<1/2$, the noncollinear spin configuration has a more complex 2:1:1 structure. The 2:2 state, which is one of the classical ground states in the entire range $0<h<h_{\rm sat}$, is stable, in the presence of a biquadratic exchange, in the low-field region only [96].
Since the plaquettes $\alpha$ share edges, one spin participates in several plaquettes. As a result, one faces the problem of extending a single-block solution to the entire lattice. This leads to additional degeneracy for the lattice spin model for each of the single-block solutions in Fig. 1. The corresponding degeneracy is most easily visualized for the uuuduuud state [72]. Specifically, consider the most symmetric uuuduuud configuration. For that we choose the down spin, for example, in the lowest left corner of the first square plaquette and extend this configuration in a translationally invariant way over the whole lattice. This state consists of parallel, vertical and horizontal, lines of up-down spins separated by lines of up spins. The alternating up-down spins along any of these lines can be freely rotated around an arbitrary angle without changing the total classical energy [72]. Since the uuuduuud state permits a maximum of such deformations, one may expect it to be selected by thermal fluctuations via the “order-by-disorder” mechanism [97, 98, 99, 100, 101, 102]. A subtle property of the order-by-disorder mechanism is a competition between a local fluctuation contribution manifested by the effective biquadratic coupling and the extended zero-energy modes related to spin distribution over the entire lattice. Furthermore, in the latter case there is the possibility of distinct effects from the thermal fluctuations, which rely on the entropic contribution of the low-energy (zero) modes, and from the quantum fluctuations, with magnons at different momenta contributing equally. As we shall see below, all these exotic features of the order-by-disorder effect are relevant for the FSAFM.
@@ II.2 Field-temperature phase diagram
Figure 2 presents the constructed $h$-$T$ phase diagram for $J_{2}/J_{1}=1/2$. First, we observe the expected stabilization of the uuuduuud phase by thermal fluctuations in the regime $T\lesssim J_{1}/10$. The stability region is asymmetric around $h_{c}=0.5h_{\rm sat}$ having a significant extent into the low-field part of the diagram. Second, there are sizable regions with the 2:1:1 and 3:1 states just below or above the uuuduuud phase, respectively. The two noncollinear states possess the longitudinal LRO in the $s^{z}$ components accompanied by the quasi-LRO in transverse spin components. Mapping this from magnetic to bosonic language [106, 107, 108], these two phases can be identified as supersolids, whereas the parent uuuduuud phase is an insulating state with the density-wave crystal.
Figure 3: The magnetization curve $m(h)$ (lines, right axis) and field derivative of the magnetization (symbols connected by lines, left axis) for the classical FSAFM with $J_{2}=0.5$, $J_{1}=1$ at three different temperatures obtained on lattices with $L=32$.
Figures 3 and 4 illustrate some of the data used in the construction of the phase diagram Fig. 2. According to the analysis at the beginning of this section, the magnetization curve $m(h)$ at zero temperature is a straight line until the saturation field $h_{\rm sat}$. Figure 3 shows some magnetization curves at low but finite temperature. The main feature that emerges at intermediate fields is a small dip for $m\lesssim 1/2$ that is most pronounced for $T/J_{1}=0.03$. We note that the improved Monte Carlo results do not reproduce the hysteresis that appeared in earlier work [72]. To better exhibit these features in the magnetization curve, we consider the field derivative of the magnetization ${\rm d}m/{\rm d}h$, also shown in Fig. 3. At zero temperature $m(h)$ has a constant slope for all fields below the saturation field $h_{\rm sat}$, i.e., ${\rm d}m/{\rm d}h=J_{1}/8$ at $T=0$ and for $h<h_{\rm sat}$. At $T>0$, we observe in Fig. 3 the emergence of small but sharp and distinct maxima in ${\rm d}m/{\rm d}h$, which may be taken as first indications of field-induced phase transitions. For $h\lesssim 4\,J_{1}=h_{\rm sat}/2$, we observe a drop in ${\rm d}m/{\rm d}h$. The slope of the magnetization curve $m(h)$ remains finite and thus one finds only a small dip but no strict plateau in the classical magnetization curve. Nevertheless, the reduced slope of the magnetization curve can be considered as a distinctive feature of the uuuduuud phase.
@@ III Quantum model
$\left.\omega(k_{x},k_{y})\right|_{J_{2}=J_{1}/2}$ $=$ $2sJ_{1}\left(\left(1+\cos k_{x}\right)\left(1+\cos k_{y}\right)-4\right)$ (8)
@@
$\left.\omega(k_{x},\pi)\right|_{J_{2}=J_{1}/2}=\left.\omega(\pi,k_{y})\right|_{J_{2}=J_{1}/2}=-8sJ_{1}+h\,$, (9)
@@ III.1.1 Magnetization curves
Figure 8 presents zero-temperature magnetization curves, where we normalized the saturation magnetization to $m=1$. In addition to finite-size results, we show estimates for the behavior of the infinite system by the dashed red lines. Following Ref. [114], this estimate is in general obtained by connecting the midpoints of the finite-size steps for the largest available system size $N$. The exception are suspected plateaux: here we use the corners of the steps instead. The main bias in this extrapolated curves is whether or not we expect a plateau and we refer to the discussion below for further details on this point.
For reference, we include the plain square-lattice case with $J_{2}=0$ (Fig. 8(a)) that has been investigated in detail previously, see for example Ref. [115], including the $N=40$ results. Here, we find a smooth curve in the thermodynamic limit with an upwards curvature that ends in a logarithmic singularity. This behavior is characteristic of a two-dimensional quantum spin system, and has been analytically predicted for the present case by second-order spin-wave theory [116].
@@
We note that a Chern-Simons theory predicted a distinct $m=1/3$ plateau in addition to the $m=1/2$ one [76]. The $N=36$ and $48$ systems indeed exhibit a broader step for $J_{2}/J_{1}=0.55$ (Fig. 8(c)) and $0.6$ (Fig. 8(d)) while the magnetization curves for the other system sizes appear to exhibit a smooth behavior in this region. The appearance of a strict $m=1/3$ plateau is in fact only possible for system sizes divisible by three. It is difficult to say on the basis of the present data if sizes $N$ that are multiplies of three favor a finite-size plateau or if sizes not commensurate with three miss a one-third plateau. We therefore chose to follow Ref. [81] and have not drawn an $m=1/3$ plateau in our extrapolated curves.
Lastly, we note that at $J_{2}/J_{1}=0.5$ and $J_{2}/J_{1}=0.55$ we have drawn an $m=0$ plateau in Fig. 8(b) and (c), respectively. This corresponds to a spin gap in zero field, as is to be expected for the paramagnetic phases around $J_{2}/J_{1}=1/2$ mentioned in the Introduction. The situation for $J_{2}/J_{1}=0.6$ (Fig. 8(d)) is less clear. Here, the $N=36$ system suggests the presence of a spin gap, but this is not confirmed by the $N=40$ system. We therefore have not drawn an $m=0$ plateau although we cannot exclude the existence of a small $m=0$ plateau, i.e., spin gap for $J_{2}/J_{1}=0.6$. In any case, we recall that the zero-field behavior is a challenging issue that is beyond the scope of the present work.
@@ III.1.2 $m=1/2$ plateau
Now we focus on a more detailed characterization of the $m=1/2$ plateau. First, we look at the width of the $m=1/2$ step, calculated as $\Delta h=E(N/2+1)-2E(N/2)+E(N/2-1)$, where $E(S^{z})$ is the energy of the ground state for the given value of $S^{z}$ at $h=0$. Figure 9 presents the behavior of $\Delta h$ as a function of $J_{2}/J_{1}$ for systems with $32\leq N\leq 52$ spins 1/2. We have chosen lattices with an aspect ratio as close as possible to unity for a given size $N$, but nevertheless the results fall into two groups, namely the square respectively rectangular geometries for $N=36$ and $48$ on the one side, and the tilted cases $N=32$, $40$, and $52$ on the other side. This non-monotonic finite-size behavior renders an extrapolation to $N\to\infty$ difficult. Nevertheless, we observe a clear and consistent enhancement of the $m=1/2$ step size in the region $0.52\lesssim J_{2}/J_{1}\lesssim 0.7$, which we thus take as a first estimate for the stability region of the $m=1/2$ plateau.
To gain further insight into the state of the $m=1/2$ plateau, Fig. 2 shows real-space spin-spin correlations at distance $R$ for the $N=52$ lattice of Fig. 7, i.e., at $S^{z}=13$. Note that $\langle s^{z}_{R}\rangle=m/2=1/4$. We therefore subtract a background of $\langle s^{z}_{0}\rangle\langle s^{z}_{R}\rangle=1/16$ from the longitudinal correlation function $\langle s^{z}_{0}s^{z}_{R}\rangle$ to obtain the connected one, that is shown in Fig. 2(a). The transverse one in Fig. 2(b) is obtained as $\langle s^{\perp}_{0}s^{\perp}_{R}\rangle=\langle\bm{s}_{0}\cdot\bm{s}_{R}\rangle-\langle s^{z}_{0}s^{z}_{R}\rangle$. Overall, we observe three distinct regimes in Fig. 2: (i) For $J_{2}/J_{1}\lesssim 0.5$ the transverse correlations dominate. This is consistent with Néel order in the transverse components, as expected in the limit $J_{2}=0$. (ii) For $0.52\lesssim J_{2}/J_{1}<0.7$, the transverse correlations essentially vanish, but strong longitudinal ones develop. This is consistent with the collinear uuuduuud state that is expected for an $m=1/2$ plateau from the investigation of the classical model in Sec. II. (iii) For $J_{2}/J_{1}>0.7$, the longitudinal components vanish again while a new signal develops in the transverse components. Different signs reflect an order that differs from the one in regime (i). Indeed, for $J_{1}=0$, one has two interpenetrating square lattices, each of which develops independent Néel order. A weak coupling $J_{1}>0$ then leads to a “stripe” phase, consistent with the correlations observed in Fig. 2(b) for $J_{2}/J_{1}>0.7$.